### 26 Feb 2024 v1

# Towards Open-ended Visual Quality Comparison

Abstract
Comparative settings (e.g. pairwise choice, listwise ranking) have been adopted by a wide range of subjective studies for image quality assessment (IQA), as it inherently standardizes the evaluation criteria across different observers and offer more clear-cut responses. In this work, we extend the edge of emerging large multi-modality models (LMMs) to further advance visual quality comparison into open-ended settings, that 1) can respond to open-range questions on quality comparison; 2) can provide detailed reasonings beyond direct answers. To this end, we propose the Co-Instruct. To train this first-of-its-kind open-source open-ended visual quality comparer, we collect the Co-Instruct-562K dataset, from two sources: (a) LMM-merged single image quality description, (b) GPT-4V “teacher” responses on unlabeled data. Furthermore, to better evaluate this setting, we propose the MICBench, the first benchmark on multi-image comparison for LMMs. We demonstrate that Co-Instruct not only achieves 30% higher superior accuracy than state-of-the-art open-source LMMs, but also outperforms GPT-4V (its teacher), on both existing related benchmarks and the proposed MICBench.
Keywords:
Large Multi-modality Models (LMM) Visual Quality Assessment Visual Quality Comparison Visual Question Answering
1 Introduction
Image quality assessment (IQA) has been an important domain in visual computing, as it provides effective recommendation [58] on high-quality visual contents and valuable guidance [54, 62, 66] for potential improvements. Most recently, several pioneer studies [51, 50, 52] have explored large multi-modality models (LMMs, e.g. GPT-4V) [34, 9, 25, 57, 4, 5], on expanding IQA from giving a scalar score (e.g. 3.457) to the open-ended scenarios, that allows evaluations in response to open-range questions, and provide detailed reasonings beyond an overall score.
While these studies sufficiently emulate human ability on IQA, they also suffer from the same drawback as human: ambiguity on absolute evaluations. For instance, as shown in Fig. 2 (a), different human observers hold different standards on the exposure level on single images, and henceforth provide diversified absolute evaluations. Nevertheless, while asked to compare the exposure level of the images, all observers agree with the rank (1)>(2)>(3)>(4) (Fig. 2(b)); all observers also agree that (1) has the highest exposure, though not all choose the option high while evaluating it independently. Given this observation, comparison has been a traditional human study setting for quality evaluation [43] and adopted by a wide range of existing subjective studies [35, 63, 10, 54, 53]. Furthermore, to avoid the ambiguity, the comparative settings are also predomintantly adopted [66, 54] while applying IQA algorithms for improvement guidance.
While comparison has widely-recognized significance for IQA, existing related datasets [35] and methods [64, 28] are generally based on overall quality comparison and have not extended to the open-ended scenarios; on the other hand, open-source LMMs [51, 25, 4] are usually only fine-tuned with single image instruction tuning datasets [26, 13, 1] and proved to lack enough capability even on two image comparison settings [68, 70]. While these gaps have clearly indicated the need of a specific instruction tuning dataset for visual quality comparison, it is too expensive to collect such a dataset from human. To avoid costly human labors, we propose an alternative strategy to collect the training dataset, named Collaborative Instruction Tuning from Weak Supervisors (Co-Instruct). Specifically, we adopt two non-perfect supervisors: 1) Merge2Compare. Originated from single image human quality descriptions of 19K images in Q-Pathway [51], we randomly match them into 100K groups with removing the most similar descriptions with an text embedding model [44]. Then, similar as the construction of LLaVA-150K [26], we prompt a single-modal LLM [16] to compare multiple human descriptions in a group, and “merge” them into 100K pseudo comparisons (as illustrated in Fig. 3(a)). 2) Teach2Compare. Observing that GPT-4V has especially high accuracy on pairwise settings [70, 68] among existing LMMs, following existing practices [2, 17], we leverage GPT-4V responses to further expand our dataset. We collect 9K unlabeled images and randomly match into 30K image groups (2-4 images per group), and obtain GPT-4V responses on both caption-like general comparisons and question-answer pairs for comparisons. By integrating Q-Instruct-200K [51] (on single images), Merge2Compare, and Teach2Compare, we construct the Co-Instruct-562K, the first instruction tuning dataset designed for open-ended multi-image quality comparison.
To correctly refer to each specific image during conversation, we define a specific image-text interleaved format [5] to handle multi-image cases, as follows:
User: The first image: <img> The second image: <img> ... <query>
Assistant: <response>
As we need to feed multiple images together during instruction tuning, adopting the most popular LLaVA [26, 39, 25] structure that linearly projects visual embeddings will exceed the context window of the language models [42] and cause errors. Henceforth, we adopt an alternative visual abstractor structure [57] to first reduce visual token length (from 1025 to 65 tokens per image), and then concatenate them with text embeddings to pass to language decoders. By learning from the Co-Instruct-562K dataset and the specially-designed input structure, we present the Co-Instruct, with up to 86% improvements than its baseline [57], and 61% better than the state-of-the-art open-source LMM. More importantly, despite using GPT-4V as one of its teachers, it still surpasses the GPT-4V teacher in a variety of multi-choice question (MCQ) benchmarks, and matches GPT-4V ability in scenarios requiring detailed language reasonings.
After training the model Co-Instruct, our another concern is the lack of abundant evaluation settings on multi-image comparison: while Q-Bench [50, 68] series have covered multiple formats on single images and image pairs, there is no existing evaluation scenario for quality comparison beyond two images. To complement this evaluation setting, we construct the MICBench. Specifically, the MICBench contains 2,000 multi-choice questions (MCQ) comparing the quality or related attributes among a group of three or four images (each half), in which over half of the questions are Which questions (e.g. which image has highest clarity?). Aiming to extract an image with a specific quality-related appearance from a group, Which questions are the most important questions related to image comparison. Despite Which questions, the MICBench also includes Yes-or-No questions and other question types (What/How/How-Many, etc) to provide a holistic benchmark setting for multi-image quality comparison.
In summary, we conduct a systematical study towards open-ended visual quality comparison. Our contributions can be summarized as three-fold:
1.
We construct the first instruction-tuning dataset for visual quality comparison, the Co-Instruct-562K. With data collected from two “weak supervisors”, Merge2Compare (LLM-merged comparisons) and Teach2Compare (GPT-4V pseudo-labeled comparisons), our public dataset significantly expands the capabilities of open-source LMMs on visual comparative settings.
2.
We propose the most capable model for open-ended visual comparison, the Co-Instruct. With image-text interleaved input format and fine-tuned with the Co-Instruct-562K dataset, it significantly outperforms existing methods (and even GPT-4V) in multiple open-ended visual comparison tasks. With open weights, it allows for broader application than proprietary models.
3.
We construct the benchmark, the MICBench, as the first benchmark to evaluate LMMs for quality comparison on multiple images (more than two). It covers 2,000 diverse-type open-range MCQs related to visual quality comparison among three or four images. The MICBench contributes to more holistic evaluation studies on the visual quality comparison problem.
2 Related Works
2.1 Visual Quality Comparison
Visual quality comparison (especially paired comparison) is a widely used subjective quality assessment methodology, serving as the most reliable way to collect human opinions [31]. However, when the number of images increases, the experiments become infeasible because of the exponential growth of pairwise comparisons [14]. While many active sampling methods have been proposed to reduce the number of pairs [55, 21, 33], they are computationally expensive and unsuitable for large-scale experiments. Despite subjective studies, learning to rank quality is widely proven as effective by many objective approaches [27, 30, 8, 64, 46, 48, 65]. Nonetheless, they typically only predict a scalar score or a binary judgement for overall comparison, limiting their ability to provide meaningful feedbacks into specific types of distortions or preferences.
2.2 LMMs for Visual Quality Evaluation
Several recent studies have explored LMMs for visual quality evaluation. The Q-Bench [50] proposes a holistic benchmark for LMMs on low-level perception (quality-related MCQs), description (quality-related captioning) and assessment (predicting scores). Following this path, Q-Instruct [51] advances the ability of LMMs with a large-scale human-annotated dataset, and Q-Align [52] designs a text-guided instruction tuning for score predictions and outperforms non-LMM approaches. However, these explorations are based on single images and have not covered comparative settings. While most recent benchmarks [68, 70] suggest that open-source LMMs trained with single image datasets cannot perform well on comparative settings, to bridge this gap, we collect the first instruction tuning dataset to teach LMMs to compare visual quality, the Co-Instruct-562K, and our model significantly improves the ability of open-source LMMs on comparative settings, moving a step forward on the basis of existing explorations.
| Subsets | Q-Instruct-200K [51] | Merge2Compare | Teach2Compare-general | Teach2Compare-Q&A | All |
| Instruction Type | Detailed Reasoning, Question Answering | Detail Reasoning | Detail Reasoning | Question Answering | – |
| # Total Images | 19K (shared, both using Q-Pathway images) | 9K (using shared images) | 28K | ||
| # Total Data Items | 202K | 100K | 30K | 230K | 562K |
[…]
| # Image Pairs | 0 | 70K | 18K | 134K | 222K |
| # Groups of Three | 0 | 20K | 6K | 51K | 77K |
| # Groups of Four | 0 | 10K | 6K | 45K | 61K |
3 Data Construction
In this section, we elaborate on the construction process of the Co-Instruct-562K dataset. Though human annotation is the most direct approach to collect data, as is widely acknowledged [21, 55], acquiring sufficient comparative data on a large set of images demands a markedly increased volume of human annotations [10, 35] in contrast to gathering opinions on the same set of individual images. To avoid this unbearable cost, we propose an alternative data construction strategy without additional human annotation, by following three key principles:
1.
Convert: Utilize reliable information from existing datasets.
2.
Learn-from-Model: Leverage the verified capabilities of models.
3.
Co-Instruct: Collecting diverse subsets that complement each other.
Under the principles, we collect two different subsets for instruction tuning: Merge2Compare (Sec. 3.1), which converts information from human quality descriptions on single images, and the ability of single-modal LLMs on comparing and analyzing texts; and Teach2Compare (Sec. 3.2), which leverages the verified ability of GPT-4V on comparing images. Finally, we discuss how the two subsets complement each other (Sec. 3.3) under a co-instruct scheme.
3.1 Merge2Compare.
In this part, we define the construction process for Merge2Compare includes three steps: 1) pair/group matching; 2) top-similarity pair removal; and 3) LLM merging. An examplar illustration of the process is shown in Fig. 4.
Step 1: Pair/Group Matching (Fig. 4(a)).
To best utilize information from existing single image quality descriptions, following the empirical rule to sample pairwise combinations to effectively rank among all individual items in a set [37, 45], we randomly sample 81K image pairs from all 19K images in the Q-Pathway. Despite pairs, we further sample 27K groups with three images and 18K groups with four images to cover the scenarios of more images.
Step 2: Top-Similarity Pair Removal (Fig. 4(b)).
The effectiveness of the merge comes from the differences among descriptions, e.g. between The quality is acceptable for <img> and The quality is poor for <img>. However, if descriptions in a pair/group contains almost the same information (e.g. both images with The clarity is good, but lighting is dark), the merged comparisons will lack enough information or even with false predictions. Henceforth, we use E5-Mistral [44] text embedding model to compute similarities among descriptions, and remove if any high-similarity description pairs exist in the group. After removal, 70K image pairs (86% of initial samples), 20K groups of three (74% of initial) and 10K groups of four (55% of initial) are preserved and fed into the LLM for merging.
Step 3: LLM Merging (Fig. 4(c)).
The key step for the Merge2Compare is to prompt LLMs to convert the single image evaluations to comparative texts. Specifically, following many existing practices [26, 69, 59], we put the descriptions as alternates of images in the context. Denote the description for image <img> as <desc>, the user query for LLM merging is formulated as follows:
(Pairs) The first image: <desc> The second image: <desc>
Which image has better quality, and why?
(Groups of Three/Four) The image: <desc>
Please rank the quality of the images and justify your rankings.
where represents the ordinal form of , e.g. is first, is second.
The merged comparisons are overall comparisons with reasonings (Fig. 4 right). To validate their reliability, we conducted a human examination on 250 random samples from Merge2Compare mixed with merged comparisons from 250 random removed groups. The correctness rate in Merge2Compare is 96%, while it is only 72% in the removed groups, demonstrating the effects of the top-similarity removal process in ensuring the quality of training data.
3.2 Teach2Compare.
Given existing evaluations [70, 68] suggesting that GPT-4V is decent at comparing visual quality (Tab. 2/3/4), we propose to collect GPT-4V responses as pseudo labels for quality comparison. As shown in Fig. 5, we collect diverse unlabeled images and feed them to GPT-4V with different prompts to obtain Teach2Compare-general (overall quality comparison) and Teach2Compare-Q&A (question-answer pairs related to quality comparison). Details as follows.
Collection of Images (Fig. 5(a)).
For Teach2Compare, we collect 9K images from various sources to cover different quality concerns and visual appearances: 1) 5.4K in-the-wild images from YFCC-100M [41] database; 2) 1.8K images artificially-distorted images from COCO [3] (with 15 types of distortions via ImageCorruptions [32]) and KADIS-700K [24] (25 types of distortions); 3) 1.8K AI-generated images from ImageRewardDB [54]. These 9K diverse unlabeled images are further grouped into 18K pairs, 6K groups of three, and 6K groups of four, for GPT-4V to provide pseudo labels under two formats, as follows.
Teach2Compare-general (Fig. 5(b)).
Similar to Merge2Compare, the general subset also consists of overall comparison with reasonings. Specifically, we substitute the <desc> in the Merge2Compare prompt template to respective real images <img> to feed to GPT-4V. After collection, we also conduct a similar 250-sample spot check on the output pseudo labels, which reports around 94% correctness. Though slightly less accurate than Merge2Compare (96%), examiners also report that GPT-4V labels contain more content information which has been observed to enhance quality understandings of models [47, 19, 20]. The two subsets are expected to complement each other for better learning outcomes.
Teach2Compare-Q&A (Fig. 5(c)).
Despite general comparisons, for GPT-4V, we also collect a specific subset to improve LMM ability on responding to open-range questions. To achieve this, we first list reference aspects (clarity, lighting, color, etc) and then ask GPT-4V to generate questions (and respective correct answers, and false answers for the questions) on comparing these aspects among the images. After removing failed generations, we obtain 230K question-answers from 30K image groups. These question-answers are converted to both direct question-answering and multi-choice questions (as A-OKVQA[38]) for training.
3.3 Rationale of Combinations.
As discussed in principle 3, our motivation is to collect subsets that can complement each other. This complementarity is reflected in the following two aspects. Firstly, in terms of general comparisons, Merge2Compare has higher accuracy but lacks fine-grained comparison (excluded by Top-similarity Pair Removal), while Teach2Compare-general, although slightly less accurate, offers more diverse scenarios and includes content information as background. Joint training of both contributes to a more comprehensive quality comparison by our model. Additionally, Teach2Compare includes a unique Q&A subset, which significantly enhances the model’s ability to answer open-range questions.
4 The Co-Instruct Model
In this section, we discuss the proposed model, Co-Instruct. Specifically, we have made two non-trivial adaptations for the multi-image comparative setting:
Visual Token Reduction (Fig. 6 (a)).
Most state-of-the-art LMMs [25, 39, 5] have adopted the simple projector that keeps a large number of tokens (e.g. 1025). This structure is not friendly to multi-image scenarios: passing only two images will exceed the max length of LLaVA [26] (2048), and four images will exceed the context window of LLaMA-2 [42] (4096). Thus, we adopt another widely-used abstractor [4, 56, 57] structure to reduce token numbers before feeding the visual embeddings to LLM, so as to easily adapt to multi-image scenarios.
Image-text Interleaved Format (Fig. 6 (c)).
Typical single-image instruction tuning usually does not care about “position of images”. Most approaches [69, 25, 4] directly pile all images before texts (<img>(<img>)<text>). While this will not cause much troubles if we query about joint information from multiple images, it is observed to introduce hallucinations across images during comparison. To solve this, we propose an image-text interleaved format for multi-image training, that each image is started with explicit text to identify its nominal:
User: The first image: <img> The second image: <img> () <query>
Assistant: <response>
In our experiments, we demonstrated that this interleaved format significantly enhances the performance of Co-Instruct (Tab. 8), notably better than using learnable special tokens (<img_st> and <img_end>) to divide images.
5 The MICBench
In this section, we discuss the proposed MICBench to cover the open-ended evaluation settings on groups of three or four images, as a complementary of existing evaluation settings (Sec. 6.3). It contains 2,000 groups of open-range questions equipped with multiple candidates, with details elaborated as follows:
Sourcing Diverse Image Groups.
To improve the diversity of the benchmark, in the MICBench, we sample image groups from two sources: (1) 400 groups of three and 400 groups of four from the images in LLVisionQA [49], which are originally sourced from 9 datasets [18, 67, 6, 12, 60, 7, 24, 15, 3]; (2) 600 groups of three and 600 groups of four on 1,000 random images sampled from unlabeled databases [41, 24, 3, 54] (zero overlap with training-set images). With in-total 2,000 groups, the MICBench contains a wide variety of quality issues and low-level appearances, providing a non-biased evaluation on quality comparison.
Evaluation Form: Multi-choice Questions (MCQs).
As the most popular evaluation form for LLM/LMM benchmarks [11, 61, 29, 49], multi-choice question (MCQ) is adopted as the evaluation form of MICBench. As is shown in Fig. 7, each image group is associated with a expert-crafted question that compare quality or related attributes among the images. Despite common question types (Yes-or-No/What/How, etc), the MICBench also introduces a special type of question, the Which questions (Fig. 7(a)), to cover this common type of human query on comparison. In total 10 human experts participate in annotating the MICBench, and the answer of each MCQ is cross-examined by another expert. Similar as existing benchmarks [29, 49], MICBench is further divided into a dev set (1,004) for method development (answers will be public), and a test set (996) to evaluate performance of LMMs (answers will be hidden from public).
6 Evaluation
6.1 Implementation Details
The Co-Instruct is fine-tuned after the released checkpoint of mPLUG-Owl2 [57], with LLaMA-2 [42] as LLM and CLIP-ViT-L14 [36] as visual embedding module. Images are padded to square and then resized to before fed into the model. The learning rate is set as , with two epochs under batch size . The final checkpoint is used for evaluation. To avoid over-fitting, only dev subsets of evaluation datasets are used to choose best training hyper-parameters, where the final reported results are from non-overlapped test subsets. All parameters are updated during training, costing in total 25 hours on 8*NVIDIA A100 GPUs.
6.2 Baseline Models
We choose five open-source recent state-of-the-art LMMs that supports multi-image inputs to compare with: LLaVA-v1.5-13B [25], InternLM-XComposer2 [5], BakLLaVA [39], EMU2-Chat [40], mPLUG-Owl2 [57] (baseline of Co-Instruct). Additionally, we also compare with three well-recognized proprietary close-source models: Qwen-VL-Max, Gemini-Pro, and GPT-4V (teacher of Co-Instruct).
6.3 Results on Existing Evaluation Settings
Despite the MICBench (Sec. 5), we also evaluate the proposed Co-Instruct against baseline models on several existing visual quality evaluation/comparison benchmarks for LMMs. The evaluation settings and results are as follows.
| Sub-categories | Question Types | Low-level Concerns | Pairwise Settings | Overall | ||||
[…]
| random guess accuracy | 50.00% | 32.03% | 33.16% | 38.95% | 41.95% | 38.69% | 43.70% | 39.82% |
[…]
| (Oct/2023) BakLLava | 60.09% | 45.42% | 50.86% | 53.09% | 58.82% | 54.52% | 55.55% | 52.75% |
| (Nov/2023) mPLUG-Owl2 (baseline of Co-Instruct) | 58.07% | 36.61% | 48.44% | 47.74% | 51.90% | 45.73% | 60.00% | 48.94% |
| (Dec/2023) Emu2-Chat | 51.94% | 29.78% | 53.84% | 42.01% | 55.71% | 46.26% | 49.09% | 47.08% |
| (Feb/2024) InternLM-XComposer2-VL | 71.81% | 58.64% | 62.28% | 65.77% | 63.67% | 64.34% | 68.00% | 65.16% |
[…]
| Gemini-Pro (Proprietary) | 65.78% | 56.61% | 56.74% | 60.42% | 60.55% | 60.46% | 60.44% | 60.46% |
| GPT-4V (Proprietary, teacher of Co-Instruct) | 79.75% | 69.49% | 84.42% | 77.32% | 79.93% | 81.00% | 68.00% | 78.07% |
| Non-expert Human | 78.11% | 77.04% | 82.33% | 78.17% | 77.22% | 80.26% | 76.39% | 80.12% |
| without Multi-image Comparative Data | 60.24% | 47.46% | 48.78% | 52.81% | 53.97% | 51.42% | 59.11% | 53.15% |
[…]
Q-Bench-A1[68] is a benchmark for visual quality comparison with 1,999 expert-crafted open-range quality-related MCQs on image pairs. In Tab. 2, we compare Co-Instruct against existing open-source and proprietary models on this benchmark. Co-Instruct shows far superior accuracy than open-source LMMs: it is 64% better than its baseline (mPLUG-Owl2), 51% better than the variant without our multi-image subsets (Merge2Compare and Teach2Compare), and also 23% better than the best of them. It also outperforms Qwen-VL-Max and Gemini-Pro by a large margin (21%/33%). Additionally, though its all MCQ training data are from GPT-4V, the student (Co-Instruct) still outperforms its teacher on this MCQ evaluation set by notable 2.7%, suggesting the effectiveness of the collaborative teaching strategy. Our model is also the only LMM that surpasses the accuracy of a non-expert human (esp. on Compare subset) in this benchmark, strongly supporting the meaningful vision of using models to relieve human labors on real-world visual quality comparisons in the future.
| Dimensions | Completeness | Precision | Relevance | Sum. | |||||||||
[…]
| (Sep/2023) LLaVA-v1.5-13B | 18.77% | 73.44% | 7.79% | 0.89 | 34.66% | 38.72% | 26.62% | 0.92 | 1.02% | 34.59% | 64.39% | 1.63 | 3.44 |
| (Oct/2023) BakLLava | 29.46% | 59.77% | 10.57% | 0.80 | 40.0% | 38.08% | 21.33% | 0.80 | 2.26% | 15.06% | 82.04% | 1.79 | 3.40 |
| (Nov/2023) mPLUG-Owl2 (baseline) | 19.43% | 65.54% | 14.45% | 0.94 | 30.94% | 43.71% | 24.63% | 0.92 | 3.79% | 26.94% | 68.28% | 1.63 | 3.50 |
| (Dec/2023) Emu2-Chat | 41.25% | 54.33% | 4.42% | 0.63 | 38.11% | 36.41% | 25.48% | 0.87 | 4.12% | 38.61% | 57.27% | 1.53 | 3.03 |
| (Feb/2024) InternLM-XComposer2-VL | 13.20% | 72.17% | 14.13% | 1.00 | 31.28% | 42.13% | 25.77% | 0.93 | 1.60% | 24.17% | 72.93% | 1.70 | 3.64 |
[…]
| Gemini-Pro (Proprietary) | 18.22% | 44.48% | 36.84% | 1.18 | 34.13% | 37.95% | 27.02% | 0.92 | 0.67% | 5.91% | 92.22% | 1.90 | 4.00 |
| GPT-4V (Proprietary, teacher of Ours) | 4.09% | 31.82% | 64.09% | 1.60 | 10.44% | 45.12% | 44.44% | 1.34 | 0.18% | 1.69% | 96.35% | 1.94 | 4.89 |
| w/o Multi-Image Comparative Data | 15.25% | 65.76% | 18.32% | 1.02 | 39.44% | 40.18% | 19.62% | 0.79 | 0.09% | 9.86% | 89.02% | 1.87 | 3.69 |
[…]
Q-Bench-A2 is a benchmark setting for general and detailed visual quality comparison with detailed reasonings on image pairs. Consisting of 499 image pairs, it employs GPT to evaluate LMM responses against the golden expert-labeled comparisons on Completeness, Precision, and Relevance. As listed in Tab. 3, the Co-Instruct can remarkably improve the Completeness (+57%) and Precision (+59%) of the comparative outputs than the w/o comparative data version, but still falls a little bit behind GPT-4V on the quantitative metrics. This might be because outputs from GPT-4V are more than twice as long as Co-Instruct outputs, while GPT evaluation is observed [22] to be in favor of longer text outputs. To validate where this potential bias occurs here, we qualitatively visualize the result of different LMMs in Fig. 8. As shown in the figure, the baseline open-source LMM even confuses the information from the two images, and Gemini-Pro makes rather poor detailed reasonings. For GPT-4V, with the longest outputs among all LMMs, even it provides an incorrect overall judgment, it still gets the highest precision score (on par with Co-Instruct which has no obvious mistakes), validating the aforementioned bias towards longer outputs.
| Dataset | CSIQ | MM21 | KADID-10k | LIVEC | KonIQ-10k | SPAQ | weighted avg. | |||||||
| Model | ||||||||||||||
| (Aug/2023) IDEFICS-Instruct-9B | 0.206 | 0.570 | 0.337 | 0.338 | 0.202 | 0.552 | 0.323 | 0.492 | 0.251 | 0.479 | 0.330 | 0.474 | 0.286 | 0.470 |
| (Sep/2023) LLaVA-v1.5-13B | 0.483 | 0.423 | 0.356 | 0.149 | 0.310 | 0.137 | 0.273 | 0.162 | 0.262 | 0.403 | 0.291 | 0.156 | 0.302 | 0.224 |
| (Oct/2023) BakLLava | 0.356 | 0.235 | 0.337 | 0.244 | 0.245 | 0.166 | 0.296 | 0.159 | 0.185 | 0.217 | 0.274 | 0.146 | 0.261 | 0.185 |
| (Nov/2023) mPLUG-Owl2 (baseline) | 0.435 | 0.627 | 0.378 | 0.306 | 0.402 | 0.443 | 0.375 | 0.441 | 0.386 | 0.417 | 0.362 | 0.356 | 0.460 | 0.397 |
| (Feb/2024) InternLM-XComposer2-VL | 0.800 | 0.527 | 0.688 | 0.377 | 0.600 | 0.552 | 0.600 | 0.516 | 0.825 | 0.581 | 0.700 | 0.755 | 0.705 | 0.567 |
[…]
| Gemini-Pro (Proprietary) | 0.672 | 0.527 | 0.604 | 0.377 | 0.790 | 0.552 | 0.650 | 0.516 | 0.652 | 0.581 | 0.671 | 0.755 | 0.678 | 0.622 |
| GPT-4V (Proprietary, teacher of Ours) | 0.778 | 0.764 | 0.792 | 0.474 | 0.763 | 0.560 | 0.837 | 0.685 | 0.835 | 0.800 | 0.871 | 0.876 | 0.823 | 0.721 |
| w/o Multi-Image Comparative Data | 0.117 | 0.650 | 0.480 | 0.392 | 0.397 | 0.466 | 0.327 | 0.432 | 0.489 | 0.512 | 0.485 | 0.397 | 0.432 | 0.449 |
[…]
2AFC-LMM[70] is a benchmark setting for general quality comparison on image pairs. It prompts LMMs to make a two-alternative forced choice (2AFC) on a pair of images. The maximum a posterior estimation is utilized to aggregate comparative preferences to single-image quality scores [43]. Then, it computes Peason’s linear correlation () between regressed scores and ground truth MOS. As shown in Tab. 4, Co-Instruct outperforms all existing models in 2AFC-LMM, including GPT-4V. Co-Instruct also shows very high consistency while swapping two images. Among all datasets, the proposed model is only inferior on the MM21 [23]. Nonetheless, we observe that the Co-Instruct has higher direct comparison accuracy than GPT-4V (Co-Instruct: 55.2%, GPT-4V: 54.4%, see supplementary for full results) on it, yet the dataset contains a large proportion of extremely similar pairs, for which Co-Instruct responds a forced choice but GPT-4V will answer “two images have similar quality” (tie), which impacts the aggregated single-image quality scores. We hope this observation can inspire further research to design better evaluation settings for fine-grained comparisons.
| Sub-categories | Question Types | Quadrants of Low-level Concerns | Overall | |||||
[…]
| Distortion | Other | |||||||
| random guess accuracy | 50.00% | 28.48% | 33.30% | 37.24% | 38.50% | 39.13% | 37.10% | 37.94% |
[…]
| (Oct/2023) BakLLava | 66.46% | 61.48% | 54.83% | 51.33% | 63.76% | 56.52% | 78.16% | 61.02% |
| (Nov/2023) mPLUG-Owl2 (baseline of Co-Instruct) | 72.26% | 55.53% | 58.64% | 52.59% | 71.36% | 58.90% | 73.00% | 62.68% |
| (Dec/2023) Emu2-Chat | 70.09% | 65.12% | 54.11% | 66.22% | 62.96% | 63.47% | 73.21% | 64.32% |
| (Feb/2024) InternLM-XComposer2-VL | 72.44% | 78.13% | 67.28% | 68.00% | 75.65% | 68.15% | 81.36% | 72.52% |
[…]
| Gemini-Pro (Proprietary) | 71.26% | 71.39% | 65.59% | 67.30% | 73.04% | 65.88% | 73.60% | 69.46% |
| GPT-4V (Proprietary, teacher of Co-Instruct) | 77.72% | 78.39% | 66.45% | 71.01% | 71.07% | 79.36% | 78.91% | 74.10% |
| Non-expert Human | 82.48% | 79.39% | 60.29% | 75.62% | 72.08% | 76.37% | 73.00% | 74.31% |
| without Multi-image Comparative Data | 79.38% | 72.23% | 67.70% | 68.71% | 72.32% | 73.97% | 83.65% | 73.38% |
[…]
Q-Bench-A1. Despite the comparative benchmarks above, we also evaluate Co-Instruct on single image MCQs from Q-Bench-A1 to verify the influences of comparative tuning on single-image quality perception. As shown in Tab. 5, Co-Instruct shows 5% improvement over the variant trained with single images only, leading GPT-4V by 4%, and marks the only LLM that surpasses non-expert human. These results have demonstrated the contribution of comparative training on general quality-related understanding of LMMs, and suggested that single-image quality evaluation does not conflict with multi-image quality comparison for LMMs and can be improved together under a unified model.
6.4 Results on MICBench
| Sub-categories | Question Types | Number of Images | Overall | |||
[…]
| #questions | 220 | 594 | 182 | 503 | 493 | 996 |
[…]
| (Oct/2023) BakLLava | 68.35% | 35.01% | 52.78% | 48.51% | 42.54% | 45.56% |
| (Nov/2023) mPLUG-Owl2 (baseline of Co-Instruct) | 62.25% | 35.70% | 53.71% | 44.19% | 45.42% | 44.80% |
| (Feb/2024) InternLM-XComposer2-VL (interpolate to 5120) | 62.95% | 47.29% | 52.02% | 55.70% | 46.51%* | 51.76% |
[…]
| Gemini-Pro (Proprietary) | 75.00% | 67.37% | 66.92% | 68.71% | 70.87% | 69.79% |
| GPT-4V (Proprietary, teacher of Co-Instruct) | 80.32% | 77.28% | 78.82% | 80.32% | 77.28% | 78.82% |
| Non-expert Human | 82.27% | 78.15% | 74.31% | 77.18% | 79.55% | 78.35% |
| without Multi-image Comparative Data | 62.72% | 37.54% | 53.30% | 45.33% | 46.65% | 45.98% |
[…]
As is shown in Tab. 6, Co-Instruct provides very competitive accuracy on open-question quality comparison among three/four images, 5.7% better than GPT-4V (best existing) and 6.4% more accurate than non-expert human; open-source LMMs even struggle to obtain 50% accuracy on this setting. Specifically, for InternLM-XComposer2-VL, without reduced tokens, it cannot directly run on MICBench-MCQ, as all groups of four exceeds its original length. To evaluate it, we interpolate its context length from 4096 to 5120. Nonetheless, we still notice a significant performance degradation on its groups-of-four answers (on interpolated positions, labeled with *), demonstrating the limitations of long visual tokens on multi-image capability. This phenomenon highlights the importance to reduce visual tokens (Sec. 4) to adapt to multi-image scenarios.
6.5 Ablation Studies
Ablation on Training Data. As the proposed dataset is composed of three parts, we discuss the effects on different subsets of data on the six evaluation scenarios in Tab. 8. Through the results, we have arrived at several important conclusions: A) Even with all information from Q-Instruct-200K, the incorporation of Merge2Compare (variant (1)) still notably enhances the capability across various settings; B) Only involving the Teach2Compare (i.e. GPT-4V labels, variant (5)) cannot outperform its teacher GPT-4V; C) instead, the superiority towards GPT-4V on question-answering benchmarks is benefited from the co-instruction on hetero-sourced training subsets (variants (6) and (7)).
| Training Subset | Evaluation Scenario | ||||||||
| Variant No. | QIn | Merge | Teach | Teach | Q-Bench-A1 | Q-Bench-A2 | Q-Bench-A1 | 2AFC-LMM | MICBench |
| Reference Results of GPT-4V | 78.07 | 4.89 | 74.10 | 0.721 | 78.82 | ||||
[…]
| (2) | ✓ | ✓ | 68.67 | 4.67 | 75.12 | 0.701 | 60.34 | ||
| (3) | ✓ | ✓ | 65.44 | 4.64 | 74.38 | 0.647 | 64.82 | ||
| (4) | ✓ | ✓ | ✓ | 69.87 | 4.82 | 76.52 | 0.749 | 66.37 | |
[…]
| (6) | ✓ | ✓ | ✓ | 80.08 | 4.68 | 75.65 | 0.728 | 81.82 | |
| (7, full) | ✓ | ✓ | ✓ | ✓ | 80.18 | 4.82 | 77.31 | 0.754 | 83.33 |
| Format | Q-Bench-A1 | Q-Bench-A2 | Q-Bench-A1 | 2AFC-LMM | MICBench |
| <img>(<img> ...) (baseline, popular strategy [4, 25, 69]) | 76.37 | 4.73 | 74.11 | 0.729 | 78.92 |
[…]
| The input image: <img> (The input image: <img>) | 78.28 | 4.80 | 75.12 | 0.749 | 82.22 |
| The first image: <img> (The second image: <img>) | 80.18 | 4.82 | 77.31 | 0.754 | 83.33 |
Besides the composition of three subsets, we also explore the effects of data scale: in Fig. 9(a), we confirm that more general comparison data contributes to an increase in detailed reasoning capabilities, and mixed data has a better effect than homogenous data at the same scale; in Fig. 9(b), we also validate that scaling up the Q&A subset helps to improve multiple MCQ metrics.
Ablation on Interleaved Format. As shown in Tab. 8, the proposed text-image interleaved format has proved non-negligible advantages than the baseline, as well as other variants without explicitly noting image orders. The results suggest the rationale of the format on multi-image comparative settings.
7 Conclusion
In this work, we investigate the open-ended visual quality comparison problem, with the aim of a model that provides answers and detailed reasonings on open-range questions that compares quality among multiple images. To achieve this, we collect the first instruction-tuning dataset to fine-tune large multi-modality models (LMMs) for comparison, the Co-Instruct-562K, with two subsets from human annotations on single images (merged by LLMs), and GPT-4V responses. With the dataset, we propose the Co-Instruct, which not only outperforms all existing LMMs (including its teacher, GPT-4V) on visual quality comparison, but also marks the first LMM with the capability to surpass human accuracy on related settings. We further construct the MICBench, the first benchmark that evaluates multi-image quality comparison for LMMs on three and four images. We expect our work to motivate future studies on visual quality comparison.

### 04 Mar 2024 v2

@@
 To this end, we propose the Co-Instruct.
-To train this first-of-its-kind open-source open-ended visual quality comparer, we collect the Co-Instruct-562K dataset, from two sources: (a) LMM-merged single image quality description, (b) GPT-4V “teacher” responses on unlabeled data.
+To train this first-of-its-kind open-source open-ended visual quality comparer, we collect the Co-Instruct-562K dataset, from two sources: (a) LLM-merged single image quality description, (b) GPT-4V “teacher” responses on unlabeled data.
 Furthermore, to better evaluate this setting, we propose the MICBench, the first benchmark on multi-image comparison for LMMs.
-We demonstrate that Co-Instruct not only achieves 30% higher superior accuracy than state-of-the-art open-source LMMs, but also outperforms GPT-4V (its teacher), on both existing related benchmarks and the proposed MICBench.
+We demonstrate that Co-Instruct not only achieves in average 30% higher accuracy than state-of-the-art open-source LMMs, but also outperforms GPT-4V (its teacher), on both existing related benchmarks and the proposed MICBench.
 Keywords:
@@
 To avoid costly human labors, we propose an alternative strategy to collect the training dataset, named Collaborative Instruction Tuning from Weak Supervisors (Co-Instruct).
-Specifically, we adopt two non-perfect supervisors: 1) Merge2Compare.
-Originated from single image human quality descriptions of 19K images in Q-Pathway [51], we randomly match them into 100K groups with removing the most similar descriptions with an text embedding model [44].
-Then, similar as the construction of LLaVA-150K [26], we prompt a single-modal LLM [16] to compare multiple human descriptions in a group, and “merge” them into 100K pseudo comparisons (as illustrated in Fig.
+Specifically, we adopt two non-perfect supervisors: 1) Merge2Compare (Fig.
 3(a)).
-2) Teach2Compare.
+Originated from single image quality descriptions on 19K images as labeled by human in the Q-Pathway [50] dataset, we randomly match them into 100K groups (2-4 images per group) with removing the most similar descriptions with an text embedding model [44].
+Then, we prompt a single-modal large language model (LLM) [16] to compare multiple human descriptions in a group, and “merge” them into 100K pseudo comparisons.
+2) Teach2Compare (Fig.
+3(b)).
 Observing that GPT-4V has especially high accuracy on pairwise settings [69, 67] among existing LMMs, following existing practices [2, 17], we leverage GPT-4V responses to further expand our dataset.
-We collect 9K unlabeled images and randomly match into 30K image groups (2-4 images per group), and obtain GPT-4V responses on both caption-like general comparisons and question-answer pairs for comparisons.
-By integrating Q-Instruct-200K [51] (on single images), Merge2Compare, and Teach2Compare, we construct the Co-Instruct-562K, the first instruction tuning dataset designed for open-ended multi-image quality comparison.
+We collect 9K unlabeled images and randomly match into 30K image groups (also 2-4 images per group), and obtain GPT-4V responses on both caption-like general comparisons and question-answer pairs for comparisons.
+By integrating Q-Instruct-200K [50] (on single images), Merge2Compare, and Teach2Compare, we construct the Co-Instruct-562K, the first instruction tuning dataset for open-ended visual quality comparison.
 To correctly refer to each specific image during conversation, we define a specific image-text interleaved format [5] to handle multi-image cases, as follows:
@@
 Assistant: <response>
-As we need to feed multiple images together during instruction tuning, adopting the most popular LLaVA [26, 39, 25] structure that linearly projects visual embeddings will exceed the context window of the language models [42] and cause errors.
-Henceforth, we adopt an alternative visual abstractor structure [57] to first reduce visual token length (from 1025 to 65 tokens per image), and then concatenate them with text embeddings to pass to language decoders.
+Moreover, as we need to feed multiple images together to the LLM decoder, adopting the most popular LLaVA [26, 39, 25] structure that linearly projects visual embeddings will exceed the context window of the language models [42] and cause errors.
+Henceforth, we adopt an alternative visual abstractor structure [56] to first reduce visual token length (from 1,025 to 65 tokens per image), and then concatenate them with text embeddings to pass to language decoders.
 By learning from the Co-Instruct-562K dataset and the specially-designed input structure, we present the Co-Instruct, with up to 86% improvements than its baseline [56], and 61% better than the state-of-the-art open-source LMM.
@@
 | Instruction Type | Detailed Reasoning, Question Answering | Detail Reasoning | Detail Reasoning | Question Answering | – |
-| # Total Images | 19K (shared, both using Q-Pathway images) | 9K (using shared images) | 28K | ||
-| # Total Data Items | 202K | 100K | 30K | 230K | 562K |
+[…]
+[…]
 […]
@@
 Most state-of-the-art LMMs [25, 39, 5] have adopted the simple projector that keeps a large number of tokens (e.g.
-1025).
-This structure is not friendly to multi-image scenarios: passing only two images will exceed the max length of LLaVA [26] (2048), and four images will exceed the context window of LLaMA-2 [42] (4096).
+1,025).
+This structure is not friendly to multi-image scenarios: passing only two images will exceed the max length of LLaVA [26] (2,048), and four images will exceed the context window of LLaMA-2 [42] (4,096).
 Thus, we adopt another widely-used abstractor [4, 55, 56] structure to reduce token numbers before feeding the visual embeddings to LLM, so as to easily adapt to multi-image scenarios.
@@
 Most approaches [68, 25, 4] directly pile all images before texts (<img>(<img>)<text>).
-While this will not cause much troubles if we query about joint information from multiple images, it is observed to introduce hallucinations across images during comparison.
+Under this piling, multiple images are not separates and LMMs might confuse the information from different images and fail to compare well (see baseline result in Fig.
+8).
 To solve this, we propose an image-text interleaved format for multi-image training, that each image is started with explicit text to identify its nominal:
@@
 Images are padded to square and then resized to before fed into the model.
-The learning rate is set as , with two epochs under batch size .
+The learning rate is set as -, with two epochs under batch size .
 The final checkpoint is used for evaluation.
@@
 3, the Co-Instruct can remarkably improve the Completeness (+57%) and Precision (+59%) of the comparative outputs than the w/o comparative data version, but still falls a little bit behind GPT-4V on the quantitative metrics.
-This might be because outputs from GPT-4V are more than twice as long as Co-Instruct outputs, while GPT evaluation is observed [22] to be in favor of longer text outputs.
-To validate where this potential bias occurs here, we qualitatively visualize the result of different LMMs in Fig.
+This might be because outputs from GPT-4V are more than twice as long as Co-Instruct outputs, while GPT evaluation used here is observed [22] to be in favor of longer text outputs.
+To further analyze this potential bias, we qualitatively visualize the result of different LMMs in Fig.
 8.
 As shown in the figure, the baseline open-source LMM even confuses the information from the two images, and Gemini-Pro makes rather poor detailed reasonings.
-For GPT-4V, with the longest outputs among all LMMs, even it provides an incorrect overall judgment, it still gets the highest precision score (on par with Co-Instruct which has no obvious mistakes), validating the aforementioned bias towards longer outputs.
+For GPT-4V, it generates the longest outputs among all LMMs, which might be the reason that it gets a relatively high precision score even its outputs are not totally correct.
+In short, the capability of Co-Instruct in reasoning-related comparisons can match that of GPT-4V, while significantly surpassing other existing LMMs.
 | Dataset | CSIQ | MM21 | KADID-10k | LIVEC | KonIQ-10k | SPAQ | weighted avg. | |||||||
@@
 […]
-| (Oct/2023) BakLLava | 68.35% | 35.01% | 52.78% | 48.51% | 42.54% | 45.56% |
+| (Oct/2023) BakLLava (length: 20482560) | 68.35% | 35.01% | 52.78% | 48.51% | 42.54%* | 45.56% |
 | (Nov/2023) mPLUG-Owl2 (baseline of Co-Instruct) | 62.25% | 35.70% | 53.71% | 44.19% | 45.42% | 44.80% |
-| (Feb/2024) InternLM-XComposer2-VL (interpolate to 5120) | 62.95% | 47.29% | 52.02% | 55.70% | 46.51%* | 51.76% |
+| (Feb/2024) InternLM-XComposer2-VL (length: 40965120) | 62.95% | 47.29% | 52.02% | 55.70% | 46.51%* | 51.76% |
 […]
@@
 6, Co-Instruct provides very competitive accuracy on open-question quality comparison among three/four images, 5.7% better than GPT-4V (best existing) and 6.4% more accurate than non-expert human; open-source LMMs even struggle to obtain 50% accuracy on this setting.
-Specifically, for InternLM-XComposer2-VL, without reduced tokens, it cannot directly run on MICBench-MCQ, as all groups of four exceeds its original length.
-To evaluate it, we interpolate its context length from 4096 to 5120.
-Nonetheless, we still notice a significant performance degradation on its groups-of-four answers (on interpolated positions, labeled with *), demonstrating the limitations of long visual tokens on multi-image capability.
-This phenomenon highlights the importance to reduce visual tokens (Sec.
+It is also noteworthy that LLaVA series and XComposer2-VL’s original context lengths are not enough for four images as they have not reduced visual token numbers, so we have to extend their context windows to evaluate them for MICBench.
+Consequentially, all these models have experienced notably worse accuracy on groups of four (on extended context length) than groups of three (within original context length), as noted in * in Tab.
+6.
+This degradation highlights the importance to reduce visual tokens (Sec.
 4) to adapt to multi-image scenarios.